The length of the rectangle is three times as large as the given width. Determine a simplified expression for the perimeter of the rectangle. The width is 2a-1
Found 2 solutions by Sir226, advanced_Learner:
Answer by Sir226(11)

3.8+2X=12.4
2X=12.4-3.8
2X/2=8.6/2
X=4.3

Answer by advanced_Learner(501)

2a-1 is width
3(2a-1)
6a-3
perimeter=12a-6+4a-2
perimeter=16a-8